OpenStudy (anonymous):

There are 10 students in the class and 10 seats in the classroom, and each student sits in the same seat during every lesson, However, the teacher insists that no student sit in their "usual" seat during an exam. How many different seating arrangements are permissible during an exam?

OpenStudy (anonymous):

there are 10!. !=factorial 10*9*8*7*6*5*4*3*2*1

OpenStudy (anonymous):

I realize this is a derangement and I'll need to use \[D _{n} = n!(1 -\frac{ 1 }{ 1! } + \frac{ 1 }{ 2! } - \frac{ 1 }{ 3! } + ...+(-1)^{n}\frac{ 1 }{ n! })\] But I don't seem to be getting the correct answer.

OpenStudy (anonymous):

Ah nevermind i did it wrong. this is a sequence

OpenStudy (anonymous):

Pretty sure I will need to use that formula

OpenStudy (anonymous):

yes you are 100% true

OpenStudy (anonymous):

what course are you taking, didn't remember this in algebra, i did arithmetic and geometry ._.

OpenStudy (anonymous):

discrete math

OpenStudy (anonymous):

10!(1 - 1/1! + 1/2! - 1/3! + 1/4! - 1/5! + 1/6! - 1/7! + 1/8! - 1/9! + (-1)^10(1/(10!)) ok this might be right now, I'm getting 1334961

OpenStudy (anonymous):

@dan815 Does this look right to you?

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

i had to get my tutor hehe

OpenStudy (anonymous):

Haha, alright thanks.